### v1

@@ Abstract
Nonlinearity is often associated with amplification, instability, or emergent behavior — where small inputs produce disproportionately large outputs. Yet in many physical systems, nonlinearities act as limiting mechanisms, suppressing the response below linear expectations, as exemplified by the breakdown of Hooke’s law at large forces. These limiting effects can become significant in the optical control of quantum materials, where large-amplitude lattice motion is essential for reaching functional states far from equilibrium. Displacive excitation of coherent phonons, perhaps the most widely used control strategy, typically assumes linearity between the driving force on the lattice and the optically excited carrier density. However, the breakdown of this assumption remains largely unexplored, and strategies are needed to overcome the resulting nonlinear bottlenecks. Here, we show that sequential optical excitation overcomes the nonlinear amplitude saturation of a key phonon mode in the van der Waals ferroelectric Td-WTe2. Using time-resolved second harmonic generation, we track the response of the interlayer sliding mode governing ferroelectricity to both single- and double-pulse excitation. With increasing optical excitation, the phonon amplitude saturates and declines — a limitation we trace to band-specific electron–phonon coupling that weakens the driving force on the nuclei. Sequential in-phase excitation of coherent sliding motion, which outlasts the electronic excitation, avoids population of counteracting electronic states. This enables higher phonon amplitudes for the same optical energy input. We exploit the enhanced lattice response to perform high-amplitude vibrational spectroscopy in the excited state, revealing a novel form of anharmonic phonon coupling that emerges exclusively under nonequilibrium conditions, and previously elusive coherent vibrational modulations of electronic states. Our results establish a more general picture of electron-phonon and phonon-phonon coupling far from equilibrium, opening new possibilities for ultrafast ferroic switching and lattice control in quantum materials.
@@
Ultrafast dynamics, nonlinear phononics, nonlinear optics, ferroelectrics
Large-amplitude coherent vibrational motion is a powerful route to accessing nonequilibrium states of matter[1, 2, 3, 4, 5, 6, 7]. In nonlinear phononics [3], intense mid-infrared fields resonantly drive infrared-active phonons to explore the potential energy surface and trigger transitions into metastable ferroelectric [8], magnetic [9], or electronic phases [10, 11]. This approach often relies on anharmonic coupling between the driven mode and Raman-active lattice distortions associated with phase transitions. Alternatively, direct excitation of Raman-active modes via displacive excitation of coherent phonons (DECP) [12, 13] exploits the coupling between transiently populated electronic states and targeted lattice modes. At the same time, decisively large vibrational amplitudes often remain elusive in DECP due to damage thresholds or anharmonic phonon decay in highly excited states [14]. For vibrational coherence outlasting electronic excitation, sequential excitation offers a promising strategy to overcome these limitations [15, 16, 17]. However, how far sequential excitation can extend the boundaries of vibrational control under intrinsic constraints remains an open question.
@@
Here, we demonstrate that splitting the energy of an optical pulse into two temporally separated excitations overcomes amplitude saturation of the polar phonon mode in a van der Waals ferroelectric by avoiding the population of counteracting electronic states, thereby enabling access to nonlinear lattice phenomena. Using time-resolved second harmonic generation (trSHG) [18, 19] and ab initio calculations, we investigate the coupling between interlayer sliding motion and ferroelectric polarization, and trace the amplitude limit of the coherent sliding phonon to the occupation of high-lying electronic states. Re-excitation of vibrational coherences following relaxation of the electronic subsystem circumvents this saturation threshold and amplifies sliding motion beyond the single-pulse limit. Tracking the enhanced lattice response reveals a strong correlation between sliding mode amplitude, damping, and frequency, which serve as direct signatures of a new form of anharmonic coupling directly mediated by the coherent nuclear motion. At highest sliding amplitudes, a modulation of the trSHG signal at twice the sliding mode frequency signals a strong, lattice-driven modulation of electronic states and nonlinear optical properties—an effect previously obscured by limited vibrational amplitudes.
@@
As a model system, we investigate WTe2, a ferroelectric van der Waals semimetal that has recently attracted attention for its giant, non-saturating magnetoresistance and ultrafast optical switching between topologically distinct states [20, 21, 22, 23, 24, 25, 26]. In its thermodynamically stable ground state, the Td polymorph (stable for temperatures $T<565\,\text{K}$), WTe2 adopts a trilayer structure that breaks inversion symmetry perpendicular to the van der Waals planes (see Fig. 1b). An uncompensated interlayer vertical charge transfer gives rise to a ferroelectric polarization oriented along the crystallographic $c$-axis, that is, normal to the layers [27, 28, 29, 30]. Ultrafast optical excitation across a broad spectral range (from terahertz to visible) depopulates antibonding orbitals between neighboring layers [21, 25, 24, 31]. This launches an interlayer sliding mode at $f_{\text{s}}=0.24\,\text{THz}$, which pushes the system toward a high-symmetry configuration on the potential energy surface restoring inversion symmetry [21, 25, 32, 33, 34, 35] (Figs. 1b and c). Interlayer sliding displacements are expected to strongly affect the material’s ferroelectric properties on ultrafast time scales [29], though the precise interplay between structural dynamics and polarization deserves further consideration.
We use symmetry considerations to investigate the connection between structural distortions and ferroelectric polarization in WTe2. The polar Td phase belongs to the non-centrosymmetric space group $Pmn2_{1}$, while the non-polar 1T’ phase crystallizes in $P2_{1}/m$. Since these two space groups are not in a group–subgroup relation, a direct transition between them cannot be achieved by a single phonon mode. Instead, we identify the true high-symmetry reference structure in the centrosymmetric $Pnma$ space group (see Fig. 1b). From this phase, two distinct lattice instabilities emerge. First, the interlayer sliding mode ($A_{1}$$\Gamma_{2}^{-}$ irreducible representation, $f_{\text{s}}=0.24$ THz), which breaks inversion symmetry, lowers the structure to $Pmn2_{1}$, and directly controls the ferroelectric polarization $P_{z}$. Second, an intralayer shear mode ($B_{1}$$\Gamma_{3}^{+}$, $f=3.4$ THz), which transforms like the $\varepsilon_{yz}$ component of the elastic tensor, lowers the symmetry to $P2_{1}/m$, and governs monoclinicity. Consequently, the Td and 1T’ structures are connected only via a two-step pathway: excitation of the interlayer sliding mode to reach the $Pnma$ reference structure, followed by intralayer shear distortion leading to 1T’. In the Td phase, both modes influence polarization dynamics through
$F=F_{Td}+\omega_{A1}^{2}(P-P_{z})^{2}+\omega_{B1}^{2}Q_{B1}^{2}+\gamma(P-P_{z})Q_{B1}^{2},$ (1)
where $F$ is the free energy and $P_{z}\sim Q_{A1}$. In other words, the interlayer sliding couples directly to $P_{z}$, while intralayer shearing couples indirectly through nonlinear phonon-phonon interactions.
Second harmonic generation, owing to its inherent sensitivity to symmetry breaking, is an ideal probe of structural degrees of freedom governing ferroelectric order [36]. Tracking the transient symmetry changes in Td-WTe2 following single-pulse optical excitation via polarization-resolved trSHG (Fig. 1a) reveals a rapid initial drop in SHG intensity within 500 fs, concomitant with prominent oscillations at the interlayer sliding mode frequency (see Figs. 1d). While all accessible tensor elements oscillate coherently at the same frequency, a strong initial suppression is observed exclusively in the vertical lobe associated with the $\chi_{zyy}$ component (compare frames at $0.5\,\text{ps}$ and $0.5\,\text{ps}$ in Fig. 1d; for further details, see Fig. S2). We speculate that this anisotropic response results from an additional structural distortion along the $c$-axis, as suggested by density functional theory calculations under optical excitation [29].
@@
To explore the limits of displacive sliding mode excitation, we perform fluence-dependent trSHG measurements (see Fig. 1e). Increasing the absorbed fluence from $0.4-6.5\,\text{mJ}\,\text{cm}^{-2}$ progressively shifts the equilibrium position $Q_{0}$ toward the centrosymmetric 1T’ configuration (Fig. 1c) and leads to a pronounced softening of the sliding mode, reflected in both long-lived signal suppression and fluence-dependent frequency shifts (Fig. 1g). Most notably, however, the sliding mode amplitude initially increases with fluence, then saturates around an absorbed fluence of $F_{\text{sat}}\approx 1.6\,\text{mJ}\,\text{cm}^{-2}$, and subsequently declines at higher excitation levels (see Fig. 1f). This behavior is consistent with observations from transient reflectivity and time-resolved X-ray diffraction measurements on Td-WTe2 [35] and the closely related compound Td-MoTe2 [37]. This amplitude saturation places a fundamental limit on the extent to which coherent sliding mode excitation can be used to control ferroelectricity.
@@
Understanding the fluence-dependent behavior of the sliding mode amplitude requires going beyond the standard DECP model, which assumes a linear relationship between excited carrier density and coherent phonon amplitude [12]. Instead, we account for the contributions of individual electronic states at different energies to the effective electron–phonon coupling at the sliding mode frequency. To this end, we perform ab initio calculations [38, 39] to determine the driving force on the sliding mode as a function of the optically-induced transient electronic temperature (see Methods Section for details).
@@
Figure 2a displays the state-resolved driving forces for the sliding mode at electronic temperatures $T_{\text{elec}}=2000\,\text{K}$ and $T_{\text{elec}}=4000\,\text{K}$, that is, above and below the saturation threshold. The contributions of individual states are weighted by their occupation, governed by changes $\Delta f_{\text{elec}}$ in the Fermi–Dirac distribution at the transient electronic temperature following optical excitation. At low excitation laser fluences, that is, low electronic temperatures, changes in occupation of electron- and hole-like states near the Fermi energy $E_{\text{F}}$ generate net positive sliding forces. With increasing fluence, higher-energy states become populated, which instead counteract sliding mode excitation.
@@
In a real-space picture [40], the generation of photo-excited electrons and holes near $E_{F}$ induces a spatial modulation of the charge density within the unit cell. Through Coulomb interactions, this transient distribution drives nuclear displacements aligned with the sliding mode pattern. In contrast, population of high-energy states produces a more homogeneous charge density, reducing net Coulomb forces on the nuclei and thereby suppressing sliding-mode-driven lattice motion (see fluence-dependent maximum sliding amplitude $Q_{\text{max}}(F)$ in Fig. 2b (violet circles)). The saturation behavior is unique to the sliding mode, whereas higher-frequency modes exhibit a continuous, monotonic increase in amplitude. Comparison between experimental data (violet triangles) and the ab initio model (violet circles) reveals close agreement in the fluence-dependence of the sliding mode amplitude (Fig. 2c). More specifically, our calculations predict a saturation of the sliding mode amplitude at an electronic temperature of $T_{\text{sat}}=3100\,\text{K}$. In experiment, this corresponds to an absorbed fluence of $F_{\text{sat}}=1.6\,\text{mJ}\,\text{cm}^{-2}$, consistent with the experimental value. Beyond this fluence, both experiment and theory show a decline in sliding mode amplitude.
@@
The amplitude limitation restricts coherent vibrational manipulation of the material’s ferroelectric [27, 20, 29] and topological [21, 24] properties to low fluences. Yet it is the high-fluence regime, where the system approaches the centrosymmetric configuration, that offers the opportunity to harness phonon coherences for coherent on–off switching of ferroelectricity and manipulation of Weyl nodes. This motivates the search for control schemes that minimize electronic excitation while maximizing achievable phonon amplitudes. Repeated in-phase excitation of coherent phonons by optical pulse trains has been shown to selectively amplify targeted vibrational modes to high amplitudes [15, 16]. A necessary condition for such a ‘cold’ driving scheme is vibrational coherence outlasting the electronic excitation—a criterion met in WTe2, as confirmed by our results and previous time-resolved studies [34, 35, 21, 25, 32, 33].
@@
We therefore aim to control the sliding mode amplitude by a double-pulse excitation scheme (Fig. 3a) [16, 17, 41, 42, 43]. To this end, we split our pump pulses in two and record trSHG traces as a function of the mutual delay $\Delta t_{\text{p-p}}$ between the two pump pulses and the double-pump-probe ($\Delta t_{\text{p-pr}}$) delay (see Fig. 3b). Strong modulations of the vibrational amplitude for different double-pulse delays $\Delta t_{\text{p-p}}$ demonstrate coherent control over the sliding mode amplitude $A_{\text{s}}$ in the excited state (see selected SHG traces in Fig. 3c), corroborated by ab initio calculations (Fig. 3d). Comparing double-pulse SHG traces for $\Delta t_{\text{p-p}}=0$, where both pump pulses act as a single excitation, and $\Delta t_{\text{p-p}}=2T_{\text{s}}$ ($T_{\text{s}}=2\pi/f_{\text{s}}$), we find that, for equal total optical energy, sequential optical excitation yields up to 50 % higher sliding amplitudes (see Fig. 3e). This demonstrates that sequential excitation enables larger vibrational amplitudes in the targeted nonequilibrium state than those attainable with single-pulse excitation at the same total fluence (see Fig. 3g).
@@
Analyzing the sliding mode amplitude after the second excitation as a function of $\Delta t_{\text{p-p}}$ reveals a strong suppression for overlapping pump pulses compared to the delay-dependent amplitude expected in coherent control experiments (see red-shaded area near $\Delta t_{\text{p-p}}=0$ in Fig. 3f). The suppression persists for inter-pulse delays up to $\sim$2 ps (compare violet trace in Fig. 3f). This timescale agrees with the carrier lifetimes determined by time- and angle-resolved photoemission spectroscopy (trARPES) [26], supporting our hypothesis that excitation of charge carriers to high energies above the Fermi level counteracts sliding mode generation. While saturable absorption in WTe2 on comparable timescales has been reported [44], the magnitude of that effect appears insufficient to explain the pronounced sliding mode saturation and, in particular, the subsequent suppression observed here.
@@
In a next step, we exploit sequential excitation to enable large-amplitude coherent vibrational spectroscopy in the excited state. In particular, double-pulse excitation not only facilitates enhanced vibrational amplitudes but also allows precise control over the amplitude while keeping other key degrees of freedom, such as electronic excitation and temperature, constant [45]. This makes it possible to isolate correlations between the coherent sliding mode amplitude on the one hand, and its damping and frequency on the other hand (see Fig. 4a) [46].
@@
Using this approach, we find a strong correlation between the sliding mode amplitude $A_{\text{s}}$ and its damping constant $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$ across the entire amplitude range (Fig. 4a), indicating that the damping is directly mediated by coherent nuclear motion. Crucially, this decay channel is absent in equilibrium crystals, where nuclear wave packets remain centered at their equilibrium positions regardless of temperature. The observed amplitude-dependent damping therefore reflects higher-order anharmonic phonon–phonon interactions [46, 47] that emerge exclusively under non-equilibrium conditions.
The most probable relaxation pathway for the sliding mode involves a Klemens-like decay of the sliding mode into two lower-energy acoustic phonons with opposite momentum [48, 49, 50] (see Fig. 4c). Due to the low energy of the sliding mode ($\hbar\omega_{\text{ph}}\approx 1\,\text{meV}$), only acoustic modes near the Brillouin zone center are energetically compatible with such a down-conversion process. Amplitude-dependent damping traces recorded at two different combined fluences support this scenario (Fig. 4b). While the two traces exhibit the same linear correlation between amplitude and damping, the slopes $(\Delta\tau/\Delta A)_{\text{s}}$ increase with increasing fluence. We attribute this to a fluence-dependent transient temperature rise $\Delta T$, which enhances the thermal population of phonons according to the Bose–Einstein distribution. As anharmonic decay is a stimulated process [51], increased occupation of the daughter phonon modes leads to an elevated decay rate of the sliding mode. For the two absorbed fluence values of $F=2.2\,\text{mJ}\,\text{cm}^{-2}$ and $F=4.1\,\text{mJ}\,\text{cm}^{-2}$, we estimate corresponding temperature increases of $\Delta T=70\,\text{K}$ or $130\,\text{K}$, respectively (see Methods Section for details).

### v2

@@ Abstract
The breakdown of Hooke’s law marks the onset of nonlinear behaviour: when displacements become large, restoring forces weaken and conventional proportionality fails. In quantum materials, intense optical excitation can drive the crystal lattice into a similar regime, where established linear relations between light, electrons, and phonons no longer hold. Sliding ferroelectrics are particularly susceptible, as controlling their polarization requires large interlayer shifts. Displacive excitation of coherent phonons, the principal mechanism for launching structural motion, typically assumes that lattice-driving forces scale linearly with the photo-excited carrier density. Whether this linearity survives at high excitation, however, remains largely unexplored, and its breakdown can fundamentally limit accessible lattice displacements. Here we show that such nonlinear limitations can be surpassed in a sliding ferroelectric by timing, rather than strengthening the optical drive. Time-resolved second-harmonic generation reveals that the interlayer sliding phonon governing ferroelectricity saturates and even diminishes under single-pulse excitation. First-principles calculations attribute this nonlinearity to band-specific electron–phonon coupling that induces competing forces on the lattice. By splitting the optical energy into two well-timed pulses that avoid populating counteracting states, we achieve markedly larger phonon amplitudes at fixed total fluence. The resulting enhanced sliding motion exposes a regime of anharmonic phonon coupling that emerges only far from equilibrium. Our findings show that nonlinear limits in driven solids can be overcome, opening new pathways for steering lattice motion in quantum materials.
@@
Ultrafast dynamics, coherent phonons, nonlinear optics, sliding ferroelectrics
Large-amplitude coherent lattice motion provides a powerful route to controlling functionality in quantum materials [weiner_femtosecond_1990, dougherty_femtosecond_1992, forst_mode-selective_2015, basini_terahertz_2024, shi_nonresonant_2024, de_la_torre_colloquium_2021, disa_engineering_2021]. Femtosecond light pulses can drive solids far from equilibrium, creating transient structural configurations with emergent ferroic [nova_metastable_2019, disa_photo-induced_2023] or electronic properties [rini_control_2007, mitrano_possible_2016, horstmann_coherent_2020, maklar_coherent_2023]. Displacive excitation of coherent phonons (DECP) [zeiger_theory_1992, stevens_coherent_2002] harnesses photo-excited carriers to launch Raman-active modes and thereby steer atomic motion. Van der Waals sliding ferroelectrics [fei_ferroelectric_2018, yang_origin_2018, yang_laser-induced_2022, ni_mechanically_2020, zhang_emerging_2025, wu_sliding_2021] are particularly promising for such vibrational control, as their polarization arises from the relative displacement of adjacent layers and remains robust down to the few-layer limit [zhang_emerging_2025, wu_sliding_2021]. Driving this sliding motion to extreme amplitudes may even enable ferroelectric switching on ultrafast timescales [yang_laser-induced_2022, yang_origin_2018]. Yet achieving such large displacements within DECP remains challenging, as electronic excitation, damage thresholds, and anharmonic decay at high fluence often limit the attainable vibrational amplitude [teitelbaum_direct_2018].
@@
Here, we show that splitting an optical pulse into two time-separated excitations overcomes amplitude saturation of the key phonon mode in a van der Waals sliding ferroelectric, enhancing control over ferroelectric polarization and enabling access to nonlinear lattice phenomena. Using time-resolved second-harmonic generation (trSHG) [shan_giant_2021, satoh_ultrafast_2007] and ab initio calculations, we investigate the coupling between interlayer sliding motion and ferroelectric polarization, and trace the nonlinear amplitude limit to the occupation of high-lying electronic states counteracting sliding phonon generation. Staggered sliding-mode stimulation avoids this compensation, restores linearity, and thus amplifies sliding motion beyond the single-pulse limit. By exploiting the enhanced lattice response, we achieve greater control over the ferroelectric polarization and uncover a strong correlation between sliding-mode amplitude, damping, and frequency, revealing a new form of anharmonic coupling mediated by coherent nuclear motion.
@@
As a model system, we investigate WTe2, a ferroelectric van der Waals semimetal that has recently attracted attention for its giant, non-saturating magnetoresistance and ultrafast optical switching between topologically distinct states [sharma_room-temperature_2019, sie_ultrafast_2019, ali_large_2014, li_evidence_2017, guan_manipulating_2021, ji_manipulation_2021, hein_mode-resolved_2020]. In its thermodynamically stable ground state, the Td polymorph of WTe2 (stable for temperatures $T<565\,\text{K}$) adopts a trilayer structure that breaks inversion symmetry perpendicular to the van der Waals planes (see calculated potential energy surface in Fig. 1a and lattice structure in Fig. 1b). An uncompensated interlayer vertical charge transfer gives rise to a ferroelectric polarization oriented along the crystallographic $c$-axis, that is, normal to the layers [fei_ferroelectric_2018, yang_origin_2018, yang_laser-induced_2022, ni_mechanically_2020]. Ultrafast optical excitation across a broad spectral range (from terahertz to visible) depopulates antibonding orbitals between neighbouring layers [sie_ultrafast_2019, ji_manipulation_2021, guan_manipulating_2021, aoki_excitation_2022]. This launches an interlayer sliding mode at $f_{\text{s}}=0.24\,\text{THz}$, which pushes the system towards a high-symmetry configuration on the potential energy surface, eventually restoring inversion symmetry [sie_ultrafast_2019, ji_manipulation_2021, he_coherent_2016, dai_ultrafast_2015, soranzio_ultrafast_2019, soranzio_strong_2022] (see Fig. 1a and displacement pattern in Fig. 1b). Such interlayer sliding motion is expected to strongly influence ferroelectric properties on ultrafast time scales, a feature shared broadly among van der Waals sliding ferroelectrics [yang_laser-induced_2022, zhang_emerging_2025]. We point out that although the sliding mode is often referred to in the literature as a ‘shear mode’, it does not break the symmetry of the elastic tensor and is therefore not a shear mode in the strict sense. Thus, it is essential to carefully distinguish between the two terms.
In a first step, we use symmetry considerations to investigate how structural distortions connect the ferroelectric and paraelectric configurations of WTe2 (see Fig. 1b). The polar Td phase belongs to the space group $Pmn2_{1}$, whereas the nonpolar, ferroelastic 1T’ phase crystallizes in the $P2_{1}/m$ structure. While the interlayer sliding mode is often discussed in the context of the $\text{Td}\rightarrow 1\text{T'}$ transition, it should be noted that the 1T’ and Td phases are not related by a group–subgroup relationship [ding_phase_2021]. Therefore, the sliding mode does not, by itself, provide a symmetry-allowed pathway between them. Instead, we identify the centrosymmetric $Pnma$ structure as the true high-symmetry reference phase, from which both the 1T’ and Td polymorphs emerge as symmetry-lowering instabilities. The two relevant distortions are the interlayer sliding mode ($\Gamma_{2}^{-}$ irreducible representation), leading to the Td phase, and an intralayer shear mode ($\Gamma_{4}^{+}$ irreducible representation), giving rise to the 1T’ phase. The structural pathway connecting the Td and 1T’ phases thus requires both interlayer sliding and intralayer shearing, with the lowest-energy route bypassing the energetically unfavourable $Pnma$ structure through the simultaneous excitation of both modes (Fig. 1a; see Methods for details on the underlying density functional theory calculations).
In order to relate the free-energy landscape $F$ to the vibrational spectrum, we perform an anharmonic phonon expansion around the minimum energy structure (Td phase), yielding
$F=F_{\mathrm{Td}}+\omega_{A_{1}}^{2}Q_{A_{1}}^{2}+\omega_{B_{1}}^{2}Q_{B_{1}}^{2}+\gamma Q_{A_{1}}Q_{B_{1}}^{2},$ (1)
where $Q_{A_{1}}\sim(P-P_{z})$ represents the amplitude of the interlayer-sliding phonon, and $Q_{B_{1}}$ the intralayer-shear phonon. The corresponding mode frequencies are $\omega_{A_{1}}/2\pi:=f_{\text{s}}=0.24\,\text{THz}$ and $\omega_{B_{1}}/2\pi=2.7\,\text{THz}$. Thus, interlayer sliding couples directly to the ferroelectric polarization $P_{z}$, whereas intralayer shearing couples only indirectly to $P_{z}$ through nonlinear phonon–phonon interactions, parametrized by the last term in Eq. 1.
@@
SHG, owing to its inherent sensitivity to inversion-symmetry breaking, is an ideal probe of structural dynamics governing ferroelectric order [denev_probing_2011]. In our trSHG experiments, we use 130-fs pump pulses ($\hbar\omega_{\text{p}}=0.82\,\text{eV}$) to trigger structural dynamics, and detect the SHG response to the probe pulses ($\hbar\omega_{\text{pr}}=1.55\,\text{eV}$) at variable time delays $\Delta t_{\text{p-pr}}$ and optical polarizations (see Methods and Supplementary Fig. S5). Tracking the transient symmetry changes in Td-WTe2 following single-pulse optical excitation reveals an initial drop in SHG intensity within 500 fs, concomitant with prominent oscillations at the interlayer-sliding-mode frequency $f_{\text{s}}$ (see Figs. 1d).
@@
Maximizing the sliding-mode amplitude provides greater leverage over the ferroelectric polarization. To explore the limits of the linear coupling between electronic excitation and the sliding-mode amplitude, we perform fluence-dependent trSHG measurements (see Fig. 1c,d). Increasing the absorbed fluence from $0.4-6.5\,\text{mJ}\,\text{cm}^{-2}$ progressively shifts the equilibrium position $Q_{0}$ towards the centrosymmetric $Pnma$ configuration (compare Fig. 1a) and leads to a pronounced softening of the sliding mode. This is reflected in a long-lived signal suppression and fluence-dependent frequency shifts, respectively (Fig. 1f). Most notably, the sliding-mode amplitude initially increases with fluence, then reaches a maximum around an absorbed fluence of $F_{\text{sat}}\approx 1.6\,\text{mJ}\,\text{cm}^{-2}$, and even declines at higher excitation levels (see red trace in Fig. 1e). This counterintuitive reduction imposes a fundamental limit on the extent to which coherent lattice motion can manipulate ferroelectricity. Together with transient reflectivity and time-resolved X-ray diffraction measurements on WTe2 [soranzio_strong_2022] and MoTe2 [fukuda_coherent_2024], our results suggest that such saturation behaviour is a general characteristic of these systems.
@@
Understanding the fluence-dependent behaviour of the sliding-mode amplitude requires going beyond the conventional, linear DECP model, which assumes a linear relationship between excited-carrier density and coherent-phonon amplitude [zeiger_theory_1992]. Instead, we account for the contributions of individual electronic states at different energies to the effective electron–phonon coupling at the sliding-mode frequency. To this end, we perform ab initio calculations [emeis_coherent_2025, pan_origin_2025] to determine the driving force on the sliding mode as a function of the optically-induced transient electronic temperature (see Methods for details).
@@
Figure 2a displays the state-resolved driving forces for the sliding mode at electronic temperatures $T_{\text{elec}}=2000\,\text{K}$ and $T_{\text{elec}}=4000\,\text{K}$, that is, above and below the amplitude-limiting threshold. The contributions of individual states are weighted by their occupation, governed by changes $\Delta f_{\text{elec}}$ in the Fermi–Dirac distribution at the transient electronic temperature following optical excitation. At low excitation laser fluence, that is, low electronic temperature, changes in occupation of electron- and hole-like states near the Fermi energy $E_{\text{F}}$ generate net positive sliding forces. With increasing fluence, higher-energy states become populated, which instead counteract sliding-mode excitation. This behaviour is not captured by linear DECP theory and requires the state-specific ab-initio approach introduced here.
@@
In a real-space picture [nicholson_beyond_2018], the generation of photo-excited electrons and holes near $E_{\text{F}}$ induces a spatial modulation of the charge density within the unit cell. Through Coulomb interactions, this transient distribution drives nuclear displacements aligned with the sliding mode pattern. In contrast, population of high-energy states produces a more homogeneous charge density, reducing net Coulomb forces on the nuclei and thereby suppressing sliding-mode-related lattice motion (see Fig. 2b). The saturation behaviour is specific to the sliding mode, whereas higher-frequency modes exhibit a continuous, monotonic increase in amplitude. Comparison between experimental data and the ab initio model reveals close agreement in the fluence-dependence of the sliding-mode amplitude (Fig. 2b). More specifically, our calculations predict a saturation of the sliding-mode amplitude at an electronic temperature of $T_{\text{sat}}=3100\,\text{K}$. In the experiment, this corresponds to an absorbed fluence of $F_{\text{sat}}=1.6\,\text{mJ}\,\text{cm}^{-2}$, consistent with the experimental value. Beyond this fluence, both experiment and theory show a decline in sliding-mode amplitude.
@@
The amplitude limitation restricts coherent vibrational manipulation of the material’s ferroelectric [fei_ferroelectric_2018, sharma_room-temperature_2019, yang_laser-induced_2022] and topological [sie_ultrafast_2019, guan_manipulating_2021] properties to low fluences. This motivates the search for control schemes that beat the limit posed by electronic excitation and reach hitherto inaccessible phonon amplitudes. Specifically, we aim to overcome the sliding-mode amplitude limitation by a double-pulse excitation scheme (Fig. 3a) [akei_role_2025, hase_femtosecond_2015, horstmann_coherent_2020, horstmann_structural_2024, maklar_coherent_2023]. Repeated in-phase excitation of coherent phonons by optical pulse trains has been shown to selectively amplify targeted vibrational modes to high amplitudes [weiner_femtosecond_1991, nelson_prospects_1991]. Such ‘cold’ driving schemes rely on vibrational coherence persisting beyond the electronic excitation. Our single-pulse excitation experiments confirm that the sliding mode in WTe2 satisfies this condition, supported by earlier studies [soranzio_ultrafast_2019, soranzio_strong_2022, sie_ultrafast_2019, ji_manipulation_2021, he_coherent_2016, dai_ultrafast_2015].
@@
We therefore split the pump pulses in two and record trSHG traces as a function of the mutual delay $\Delta t_{\text{p-p}}$ between the two pump pulses and the double-pump-probe delay $\Delta t_{\text{p-pr}}$ (see Fig. 3b). Strong modulations of the vibrational amplitude for different double-pulse delays $\Delta t_{\text{p-p}}$ demonstrate coherent control over the sliding-mode amplitude $A_{\text{s}}$ in the excited state (see selected SHG traces in Fig. 3c), corroborated by ab initio calculations (Fig. 3d). Comparing double-pulse SHG traces for $\Delta t_{\text{p-p}}=0$, where both pump pulses act as a single excitation, and $\Delta t_{\text{p-p}}=2T_{\text{s}}$ ($T_{\text{s}}=2\pi/f_{\text{s}}$), we find that, for equal total optical energy, sequential optical excitation yields up to 50 % higher sliding amplitudes (see Fig. 3e). This key result is visualized in Fig. 3g, which directly compares the yields of single- and double-pulse excitation and shows that sequential excitation extends the linear relation between optical energy and sliding amplitude in the targeted nonequilibrium state.
@@
Analyzing the sliding-mode amplitude after the second excitation as a function of $\Delta t_{\text{p-p}}$ further reveals a strong suppression for overlapping pump pulses compared to the delay-dependent amplitude expected in coherent control experiments (see red-shaded area near $\Delta t_{\text{p-p}}=0$ in Fig. 3f). The suppression persists for inter-pulse delays up to $\sim$2 ps (compare violet trace in Fig. 3f). This timescale agrees with the carrier lifetimes determined by time- and angle-resolved photoemission spectroscopy [hein_mode-resolved_2020], supporting our hypothesis that excitation of charge carriers to high energies above the Fermi level counteracts sliding-mode generation. We note that saturable absorption in WTe2 on a few-ps timescale [gao_broadband_2018] cannot account for this behaviour as the magnitude of that effect appears insufficient to explain the pronounced sliding mode saturation and, in particular, its subsequent decline.
@@
In a next step, we exploit the fact that we can beat amplitude saturation by well-timed sequential excitation to enable large-amplitude coherent vibrational spectroscopy in the excited state. In particular, double-pulse excitation not only facilitates enhanced vibrational amplitudes but also allows precise control over the amplitude while keeping other key degrees of freedom, such as electronic excitation and temperature, constant [scharf_coherent_2025, ravnik_time-domain_2021, cheng_coherent_2017]. This makes it possible to isolate correlations between the coherent sliding-mode amplitude on the one hand, and its damping and frequency on the other hand (see Fig. 4a) [cheng_coherent_2017].
@@
Applying this scheme, we observe a strong correlation between the sliding mode’s coherent amplitude $A_{\text{s}}$ and its damping constant $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$ across the full amplitude range (Fig. 4a,b). This behaviour indicates that the damping is directly mediated by coherent nuclear motion. Crucially, such a decay channel is absent in equilibrium crystals, where nuclear wave packets remain centered at their equilibrium positions irrespective of temperature. Amplitude-dependent damping traces recorded at different combined fluences reveal a strikingly linear relation, characterized by the slope $(\Delta\gamma/\Delta A)_{\text{s}}$, with larger coherent amplitudes producing proportionally stronger damping (Fig. 4b). Notably, $(\Delta\gamma/\Delta A)_{\text{s}}$ itself increases with combined fluence, such that for a given initial sliding amplitude, higher overall excitation enhances the damping.
To identify the microscopic origin of this amplitude- and fluence-dependent damping, we consider a Klemens process  [dyson_role_2013, klemens_anharmonic_1966, hejda_hot-electron_1993] as the most probable decay mechanism for the sliding mode. As the lowest-energy optical phonon, the sliding mode can decay into two lower-energy acoustic phonons with opposite momentum (Fig. 4c). Stimulated phonon decay has been predicted to strongly enhance optical-phonon damping under nonequilibrium conditions, with the decay rate of the optical phonon proportional to the populations of the optical and acoustic modes. [romanov_effects_1999]
Consistent with this picture, the two experimental trends established above naturally emerge from stimulated decay of the sliding mode into acoustic modes. First, the damping increases with the sliding-mode population, or equivalently its coherent amplitude. Second, the damping grows with combined fluence, as higher overall excitation enhances scattering into low-energy acoustic branches. Thus, stimulated phonon decay provides a microscopic rationale for the observed amplitude- and fluence-dependent damping.